OpenStudy (anonymous):

Fan & medal Give examples of 3 different Polynomial Identities.

OpenStudy (brittanydosey):

Example 1: If x = 10, y = 5a (10 + 5a)2 = 102 + 2·10·5a + (5a)2 = 100 + 100a + 25a2 Example 2: if x = 10 and y is 4 (10 - 4)2 = 102 - 2·10·4 + 42 = 100 - 80 + 16 = 36 The opposite is also true: 25 + 20a + 4a2 = 52 + 2·2·5 + (2a)2 = (5 + 2a)2

OpenStudy (anonymous):

Thanks, this really helps! :D If you don't mind me asking, what is a Polynomial Identities?

OpenStudy (brittanydosey):

A polynomial identity is just a true equation, often generalized so that it can apply to more than one situation. For instance, (x2 + y2)2 = (x2 – y2)2 + (2xy)2 is true for any x and y.

OpenStudy (anonymous):

Ohhh, okay thanks. This helps me a lot.

OpenStudy (brittanydosey):

your welcome, anytime! ^.^

OpenStudy (anonymous):

Would you be able to help with one more question? It's related to this question.

OpenStudy (brittanydosey):

yes

OpenStudy (anonymous):

Okay, thanks. This is the question. 3) Pick one of the polynomial identities and prove that it is true

OpenStudy (brittanydosey):

this one has got me stuck im sorry my mind is not working today.

OpenStudy (anonymous):

Okay, thanks anyway.